OpenStudy (anonymous):

how would i solve this problem -6=(x/2)^3+2

OpenStudy (anonymous):

You need to isolate x. To do this you need to start using operations to get x by itself.

OpenStudy (anonymous):

how would i get rid of (x/2)^3

OpenStudy (anonymous):

Take the cubed root of both sides.

OpenStudy (anonymous):

so it would be -2=x/2

OpenStudy (anonymous):

convert it to: x^3=-64

OpenStudy (anonymous):

x^3=(-4)^3 x=-4

OpenStudy (anonymous):

where did u get the -64 ?

OpenStudy (radar):

\[(x/2)^{3}=-8\]\[(x ^{3}/-8)=-8\]\[x ^{3}=64\] obtained by multiplying both sides by -8. x=4 obtained by taking the cube root of both sides.

OpenStudy (radar):

oops a mistake i see. \[x ^{3}/8=-8\] now multply by 8 \[x ^{3}=-64\] x=-4

OpenStudy (radar):

more better lol

OpenStudy (anonymous):

(x/2)^3 is x^3/8 so the cube stays with x

OpenStudy (radar):

yes

OpenStudy (radar):

Any questions?

OpenStudy (anonymous):

why does the cube stay ?

OpenStudy (radar):

The exponent "3" applies to everything within the parenthesis both the numerator and the denominator. so both the 2 and the x are cubed.

OpenStudy (radar):

that would result in x^3/8

OpenStudy (radar):

\[(2/4)^{3}=8/64=1/8\]

OpenStudy (radar):

\[(a/b)^{3}=a ^{3}/b ^{3}\]

OpenStudy (radar):

Do you follow with understanding?